X-rays usually have wavelengths of one to ten angstroms. Calculate the energy in joules of photons with a 2-angstrom wavelength. Calculate the energy in kilojoules per mole of such two angstrom wavelength photons. Compare it with the bond energy of a carbon-carbon single bond (347 kJ/mol). Would you expect x rays to be able to participate in bond breaking chemical reactions? Why?
Answer #1

E=hc/λ

E= (6.626×10−34×3.00×108 )/2×10−10 J

E=9.9×10−16 J

1kJ = 1000 J and 1 mole = 6.023 * 1023 photons or molecule

E = (9.9×10−16 * 6.023 * 1023 ) / 1000 = 596277 kJ /moL

as we can see this energy is way higher than C-C bond energy 347 kJ/mol hence can be used to break bond in chemical reaction but since energy is very high hence it is not desirable to break bond using xray